# A Survey of Protocol Fuzzing

### v2

@@ 4.1.1. Top-Down Approaches
Top-down approaches construct the communication model of the protocol by learning from the textual description of the protocol, such as the specifications or documents. Top-down approaches require protocol specifications as input, thus mostly used by open protocols. Benefiting from global protocol knowledge of the specifications as well as the precisely defined states and transitions, communication models constructed by top-down approaches are relatively complete and accurate. It is worth noting that the constructed communication model may still differ from the implementation’s. This is because developers may customize or extend the design described in the specification according to the practical situation. Such a difference may affect the final fuzzing performance. Methodologically, there are manual and automatic ways to construct a communication model from protocol documents.
@@
Manual Construction (labeled as “manual” in Table 3 column 4): Most existing works construct a communication model manually with considerable domain expertise (Beurdouche et al., 2015; Jero et al., 2015; Somorovsky, 2016; Kim et al., 2017; Jero et al., 2018; Shi et al., 2019; Garbelini et al., 2020; Ruge et al., 2020; Peterson et al., 2020; Zou et al., 2021; Garbelini et al., 2022c; Park et al., 2022; Romero and Rivas, 2019; He et al., 2022; Amusuo et al., 2023; Ammann et al., 2024). For example, Garbelini et al. (Garbelini et al., 2020) and GREYHOUND (Garbelini et al., 2022c) construct a holistic state machine of Bluetooth Low Energy (BLE) and Wi-Fi by referring to the core design of protocol specifications (802, 2021; Chen et al., 2005; 802, 2020; SIG, 2016) to guide fuzzing. Though manually constructing a state machine is an error-prone and labor-intensive task, the benefit here is that human experts can flexibly customize (tailor or extend) the state machine to maximize its effect toward the work’s goal. For example, to detect the state machine bugs caused by incorrect multiplexing between different protocol modes, e.g., different protocol versions, extensions, authentication modes, or key exchange methods, Beurdouche et al. (Beurdouche et al., 2015) manually construct a composite state machine including all valid state transitions across protocol modes. That composite state machine is then used to generate deviant traces as testcases to discover invalid state transitions. Similarly, FuzzBD (Kim et al., 2023) is meticulously designed to accommodate the unique dual-role characteristic inherent in the USB Power Delivery (USBPD) protocol, where each device simultaneously functions as both a power source and a power sink. By integrating the state machines of these two roles, FuzzBD is capable of supporting seamless on-the-fly power role switching during the fuzzing process. Different from the above-discussed works, some works choose to learn partial information of a state machine for guidance. Zou et al. propose TCP-Fuzz, a novel approach that incorporates 15 dependency rules manually extracted from RFC documents. These rules encompass various dependencies, including packet-to-packet, syscall-to-packet, and syscall-to-syscall interactions. Utilizing these rules, TCP-Fuzz adeptly generates testcases by concurrently producing the interdependent packets and syscalls. A further exemplar is L2Fuzz (Park et al., 2022). The authors construct a map delineating the valid commands pertinent to each of the 19 states identified in the protocol. This mapping facilitates the generation of testcases that are specifically tailored to produce commands acceptable in the current state, thereby enhancing the relevance and effectiveness of the testing process. There are also some works addressing the problem that the communication models between the specification and implementation are not completely equal. Using heterogeneous Single-Sign-On (SSO) platforms as an example, MoSSOT (Shi et al., 2019) constructs a state machine of a regular SSO process first, then analyzes the practical SSO network traffics of different SSO platforms to learn the implementation details such as key parameters in each action. These implementation details refine the state transition conditions in the state machines of different SSO platforms.
@@
Automatic Construction: To automate the error-prone and labor-intensive process of manual communication model construction, some works automatically retrieve semantic constraint from protocol specification (Atlidakis et al., 2019; Pacheco et al., 2022). For example, RESTler (Atlidakis et al., 2019) learns the message dependency relationships based on the return types from Swagger specification, which is a structural specification format describing the RESTful API endpoints, methods, parameters and return types. Pacheco et al. propose to use Natural Language Processing (NLP) to extract a finite state machine (FSM) from the protocol specifications (Pacheco et al., 2022). Note that these two papers are not listed in Table 3 since these works are not building stateful protocol fuzzers. Another work that uses automatic state machine construction is CHATAFL (Meng et al., 2024), which leverage the emerging technology large language model to infer the current state of the target and generating suitable state transfer packets.
@@ 4.1.2. Bottom-Up Approaches
Bottom-Up approaches provide another solution for communication model reconstruction. These approaches utilize the observable information of a protocol implementation to reconstruct the communication model. Since they do not rely on textual documentation or specifications, they are suitable for proprietary protocols. Different from top-down approaches which have clear definitions of protocol states in the documents, the definitions of states in bottom-up approaches are purpose-specific and may vary among use cases, methods, and implementations. For example, AFLNET (Pham et al., 2020) determines a protocol state according to the status code of the PUT’s response. Another example is StateAFL (Natella, 2022), which clusters the memory layout of long-lived memory as different states. From the learning source’s point of view, these methods can be divided into two categories, namely traffic-analysis-based approaches and program-analysis-assisted approaches.
@@
Traffic-Analysis-Based approaches: The traffic-analysis-based methods focus on reconstructing the protocol communication model purely from the observed network traffic traces (de Ruiter and Poll, 2015; Zhao et al., 2019; Luo et al., 2019; Fiterau-Brostean et al., 2020; Pham et al., 2020; Wang et al., 2021). This kind of approach is easy to operate and works well in cases that the program execution cannot be traced, e.g., cannot obtain the firmware containing the target program. The traffic-analysis-based communication model construction approaches can be passive or active.
@@
Passive learning (labeled as “TAPL” in Table 3 column 4) methods mainly rely on a set of pre-collected network traces of the PUT with other entities to infer the communication model (Gascon et al., 2015; Fan and Chang, 2018; Zhao et al., 2019; Yu et al., 2022). The learning algorithms proposed by existing works can be divided into two categories: statistics-based and neural-network-based algorithms. For the former category, Pulsar (Gascon et al., 2015) builds a second-order Markov Model by computing the probability of the occurrence of the adjacent messages in the network trace corpus and then minimizes this Markov model into a DFA. After a message has been received, Pulsar matches it to one of the states in the inferred DFA to select a valid response template for building a new testcase. For the latter category, Fan et al. (Fan and Chang, 2018) and SeqFuzzer(Zhao et al., 2019) use LSTM to learn the grammar and temporal features of stateful protocols. Specifically, they employ LSTM as the encoder and decoder of the sequence-to-sequence (seq2seq) model (Sutskever et al., 2014). Seq2seq model is an encoder-decoder model structure that can handle input and output sequences of different lengths. The encoder LSTM model learns the features of the protocol via captured network traces, while the decoder LSTM model is used to generate fuzzing inputs. Passive network trace-based state machine learning methods are easy to operate and fast-running. However, the quality of the constructed state machine depends on the coverage of captured traffics. In practice, it is hard to capture a comprehensive set of message types and sequences, causing the constructed communication model lacking part of uncaptured states or state transitions.
@@
Active learning (labeled as “TAAL” in Table 3 column 4) methods involve learning the communication model during the fuzzing process. These approaches can be categorized based on whether the global state set is predefined. The first category does not define the global state set in advance, i.e., meaning it does not predetermine the number and nature of possible states in the state machine. This approach employs automata active learning algorithms to discern the state machine of the target. The learning algorithms rely on user-defined input/output alphabets and mappers between alphabets and concrete messages. Starting from an empty state machine, these algorithms iteratively propose and refine the model by interacting with the target protocol implementation, ceasing only when no counterexamples to the learned state machine are found. Most works in this category (de Ruiter and Poll, 2015; McMahon Stone et al., 2018; Daniel et al., 2018; Fiterau-Brostean et al., 2020; Aichernig et al., 2021; Fiterau-Brostean et al., 2022; Sagonas and Typaldos, 2023) utilize Angluin’s algorithm, defining input alphabets based on protocol specifications and translating these into actual messages using message templates. Conversely, the second category predefines the complete state set to circumvent the complexities of automata learning algorithms. Typically, this approach defines the state set through a rule-based method, learning transitions between states by mutating known messages. In essence, the potential states (nodes) of the state machine are predetermined, and the focus is on discovering and incorporating state transitions (edges). For example, AFLNET (Pham et al., 2020) uses response message status codes to determine the current protocol state, mutating real message sequences to uncover transitions. Bleem (Luo et al., 2023) utilizes the Scapy library to parse messages and abstracts them into various message types by retaining all fields of the enumeration type. This strategy is based on empirical observations from over 50 protocols supported by Scapy, where different enumeration field values typically signify distinct packet or frame types. Bleem then uses these abstracted message traces to construct a guiding graph for fuzzing. Another instance is Braktooth (Garbelini et al., 2022a), which defines eight rules mapping messages to states based on message characteristics. It operates as a proxy between the PUT and a standard protocol stack, mutating communications to explore additional state transitions. Similarly, Garbelini et al.(Garbelini et al., 2022b) establish mapping rules to identify states and learn the state machine using capture traces (i.e., pcap files).
@@
Program-Analysis-Assisted Approaches (labeled as “PAL” in Table 3 column 4): Compared with traffic-analysis-based approaches, program-analysis-assisted approaches additionally use internal execution information to construct the communication model. In general, the internal execution information includes the results of static and dynamic program analysis, which requires not only the access to the program but also the availability of analysis frameworks such as program instrumentation tools. Based on the type of the used internal execution information, existing works can be divided into execution-trace-based and state-variable-based. Execution-Trace-Based approaches recognize different internal execution states according to the execution trace of the target. For example, ICSFuzzer(Fang et al., 2021) dynamically instruments the target supervisory software to collect the trace. By comparing the identity of execution traces, ICSFuzzer can distinguish whether the PUT is in a different state. The state-variable-based approaches detect protocol state transitions by tracking the value changes of state variables during input processing (Natella, 2022; Stone et al., 2021; Ba et al., 2022; Qin et al., 2023). These approaches are based on a simple observation that most protocol implementations use certain variables to store the current state. Therefore, they identify these variables as state variables and use their values to distinguish different states. For example, StateAFL (Natella, 2022) identifies possible state variables by identifying long-lived data structures in memory snapshots. Similarly, STATEINSPECTOR (Stone et al., 2021) identifies state variables by locating memory regions in heap memory that kept the same values in the execution of each message sequence. Differently, SGFuzz (Ba et al., 2022) identifies state variables through regular expressions by automatically extracting all enum type variables that are assigned at least once. The insight behind this approach is based on the investigation that most protocol implementations use enum-type state variables.
@@ 4.2. Task Scheduling
In the realm of recent protocol fuzzing research, the scheduling phase has been distinctly categorized based on the methodology employed for handling state-related complexities. This classification leads to two primary categories: Hierarchical Approaches and Monolithic Approaches. Hierarchical approaches decompose the scheduling process into two discrete phases: intra-state scheduling, which focuses on testing within a single state, and inter-state scheduling, which manages the transition between states. These phases are implemented separately, allowing for nuanced control over the fuzzing process. In contrast, monolithic approaches employ a single, unified scheduling phase, which integrates state-related information as a coefficient within the scheduling algorithm. Besides state-related information, many works utilize other categories of information for scheduling purposes. However, the scheduling algorithms based on these categories of information are generally universal and have been well-discussed in the literature (Manes et al., 2021; Zhu et al., 2022b). Therefore, we did not delve into detailed discussions about them.
@@
Hierarchical Approaches. In this paradigm, the scheduler employs a two-step algorithm: (1) selection of a state for fuzzing using a state scheduling algorithm, followed by (2) application of a general scheduling algorithm to optimize fuzzing within that state. The heuristics used by the scheduling process mainly fall into three categories, namely rarity-preferred (SRHS in Table 3 column 5), performance-preferred (SPHS in Table 3 column 5), and complexity-preferred (SCHS in Table 3 column 5), as detailed in Table 4. Rarity-preferred heuristics allocate more resources to seldomly exercised states, hypothesizing that these states harbor more undiscovered adjacent states or code logics (Pham et al., 2020; Natella, 2022; Sun et al., 2019; Borcherding et al., 2023). Performance-preferred heuristics prioritize states demonstrating higher code coverage or bug discovery rates (Chen et al., 2019; Pham et al., 2020; Natella, 2022; Garbelini et al., 2020, 2022c). Additionally, some works utilize complexity-preferred heuristics, favoring states with greater complexity (i.e., connected to more basic blocks) or deeper states (i.e., further from the initial state) (Gascon et al., 2015; Fang et al., 2021). For example, ICSFuzzer (Fang et al., 2021) inclines to choose the deeper states and those states that exercise more basic blocks. As a generation-based fuzzer, Pulsar (Gascon et al., 2015) calculates the weight of all states that are reachable from the current state, and then selects the state that has the max weight to be tested next. In detail, the weight of a state is calculated as the sum of all mutable fields in a fixed number of transitions. However, as all these state selection algorithms are implemented and evaluated separately on different platforms and targets, it is hard to make a fair comparison and achieve conclusive findings. Liu et al. (Liu et al., 2021) evaluate the three existing state selection algorithms of AFLNet (Pham et al., 2020) including a rarity-preferred algorithm, an algorithm that randomly selects states, and a sequential state selection algorithm. They find that these algorithms achieved very similar results in terms of code coverage. They attribute the reasons to the coarse-grained state abstraction of AFLNET and the inaccurate estimation of the state productivity. Therefore, they propose the AFLNETLEGION algorithm (Liu et al., 2021) to address these issues, which is based on a variant of the Monte Carlo tree search algorithm (Liu et al., 2020).
Monolithic Approaches. In a monolithic manner, state-related information is calculated as a coefficient of the original scheduling algorithm used in the seed scheduling. For example, SGFuzz (Ba et al., 2022) divides the states into rare states and normal states according to the exercised times. When assigning energy to seeds, it calculates the proportion of the rare states that are exercised by each seed and adds this proportion as one of the parameters on the basis of the original power scheduling algorithm. In a similar way, SGFuzz assigns more energy to the seeds containing state transitions that correspond to the expected protocol behaviors. This is because SGFuzz expects that these valid state transitions are easier to be mutated to other invalid state transitions, thus incurring error handling logic. Similarly, LTL-Fuzzer (Meng et al., 2021) also schedules the entire seed. It prioritizes the seeds that are closer to the target code locations during the execution.

### v3

@@ 4.1.1. Top-Down Approaches
Top-down approaches construct the protocol communication model by learning from the textual description of the protocol, such as specifications or documents. Top-down approaches require protocol specifications as input, thus are mostly used by open protocols. Benefiting from global protocol knowledge of the specifications and the precisely defined states and transitions, communication models constructed by top-down approaches are relatively complete and accurate. It is worth noting that the constructed communication model may still differ from the implementation’s. This is because developers may customize or extend the design described in the specification according to the practical situation. This difference may affect the final fuzzing performance. Methodologically, there are manual and automatic ways to construct a communication model from the protocol documents.
@@
Manual Construction (labeled as “manual” in Table LABEL:table-inputgen column 4): Most existing works construct a communication model manually with considerable domain expertise (Beurdouche et al., 2015; Jero et al., 2015; Somorovsky, 2016; Kim et al., 2017; Jero et al., 2018; Shi et al., 2019; Garbelini et al., 2020; Ruge et al., [n. d.]; Peterson et al., 2020; Zou et al., [n. d.]; Garbelini et al., 2022c; Park et al., 2022; Romero and Rivas, 2019; He et al., 2022; Amusuo et al., 2023; Ammann et al., 2024). For example, Garbelini et al. (Garbelini et al., 2020) and GREYHOUND (Garbelini et al., 2022c) construct a holistic state machine of Bluetooth Low Energy (BLE) and Wi-Fi by referring to the core design of the protocol specifications (802, 2021; Chen et al., 2005; 802, 2020; SIG, 2016) to guide fuzzing. Though manually constructing a state machine is an error-prone and labor-intensive task, the benefit here is that human experts can flexibly customize (tailor or extend) the state machine to maximize its effect toward the work’s goal. For example, to detect state machine bugs caused by incorrect multiplexing between different protocol modes, e.g., different protocol versions, extensions, authentication modes, or key-exchange methods, Beurdouche et al. (Beurdouche et al., 2015) manually construct a composite state machine including all valid state transitions across protocol modes. That composite state machine is then used to generate deviant traces as testcases to discover invalid state transitions. Similarly, FuzzPD (Kim et al., 2023) is meticulously designed to accommodate the unique dual-role characteristic inherent in the USB power delivery protocol (USBPD), where each device simultaneously functions as both a power source and a power sink. By integrating the state machines of these two roles, FuzzBD is able to support seamless switching of power roles on the fly during the fuzzing process. Unlike the above-discussed works, some work chooses to learn partial information of a state machine for guidance. Zou et al. propose TCP-Fuzz, a novel approach that incorporates 15 dependency rules manually extracted from RFC documents. These rules encompass various dependencies, including packet-to-packet, syscall-to-packet, and syscall-to-syscall interactions. Using these rules, TCP-Fuzz adeptly generates testcases by simultaneously producing the interdependent packets and syscalls. Another example is L2Fuzz (Park et al., 2022). The authors construct a map delineating the valid commands pertinent to each of the 19 states identified in the protocol. This mapping facilitates the generation of testcases that are specifically tailored to produce commands acceptable in the current state, thereby enhancing the relevance and effectiveness of the testing process. Some works also address the problem that the communication models between the specification and implementation are not completely equal. Using heterogeneous Single-Sign-On (SSO) platforms as an example, MoSSOT (Shi et al., 2019) constructs a state machine of a regular SSO process first and then analyzes the practical SSO network traffics of different SSO platforms to learn the implementation details, such as key parameters in each action. These implementation details refine the state transition conditions in the state machines of different SSO platforms.
@@
Automatic Construction: To automate the error-prone and labor-intensive process of manual communication model construction, some works automatically retrieve the semantic constraint from the protocol specification (Atlidakis et al., 2019; Pacheco et al., 2022). For example, RESTler (Atlidakis et al., 2019) learns the message dependency relationships based on the return types from Swagger specification, which is a structural specification format describing the RESTful API endpoints, methods, parameters and return types. Pacheco et al. propose to use Natural Language Processing (NLP) to extract a finite state machine (FSM) from the protocol specifications (Pacheco et al., 2022). Note that these two papers are not listed in Table LABEL:table-inputgen since these works are not building stateful protocol fuzzers. With the advancement of large language model (LLM) technology, numerous studies have begun to employ LLM techniques to automate the learning and comprehension of protocol specifications, such as mGPTFuzz(Ma et al., 2024), CHATAFL(Meng et al., 2024) and LLMIF(Wang et al., 2024). These approaches aim to provide guidance during fuzzing processes, particularly in inferring the current protocol state and generating appropriate test messages. The integration of LLMs offers a promising avenue to improve the efficiency and effectiveness of protocol fuzzing by leveraging their sophisticated language understanding capabilities to interpret complex protocol specifications and guide the testing strategy accordingly.
@@ 4.1.2. Bottom-Up Approaches
Bottom-Up approaches provide another solution for communication model reconstruction. These approaches utilize the observable information of a protocol implementation to reconstruct the communication model. Since they do not rely on textual documentation or specifications, they are suitable for proprietary protocols. Unlike top-down approaches, which have clear definitions of protocol states in the documents, the definitions of states in bottom-up approaches are purpose-specific and may vary among use cases, methods, and implementations. For example, AFLNET (Pham et al., 2020) determines a protocol state according to the status code of the PUT’s response. Another example is StateAFL (Natella, 2022), which groups the memory layout of long-lived memory as different states. From the learning source’s point of view, these methods can be divided into two categories, namely traffic analysis-based approaches and program analysis-assisted approaches.
@@
Traffic-Analysis-Based Approaches: The traffic-analysis-based methods focus on reconstructing the protocol communication model purely from the observed network traffic traces (de Ruiter and Poll, 2015; Zhao et al., 2019; Luo et al., 2019; Fiterau-Brostean et al., 2020; Pham et al., 2020; Wang et al., 2021). This kind of approach is easy to operate and works well in cases that the program execution cannot be traced, e.g., cannot obtain the firmware containing the target program. The traffic analysis-based communication model construction approaches can be passive or active.
@@
Passive learning (labeled as “TAPL” in Table LABEL:table-inputgen column 4) methods mainly rely on a set of pre-collected network traces of the PUT with other entities to infer the communication model (Gascon et al., 2015; Fan and Chang, 2018; Zhao et al., 2019; Yu et al., 2022). The learning algorithms proposed by existing works can be divided into two categories: statistics-based and neural network-based algorithms. For the former category, Pulsar (Gascon et al., 2015) builds a second-order Markov model by computing the probability of the occurrence of the adjacent messages in the network trace corpus and then minimizes this Markov model into a DFA. After receiving a message, Pulsar matches it with one of the states in the inferred DFA to select a valid response template for building a new testcase. For the latter category, Fan et al. (Fan and Chang, 2018) and SeqFuzzer(Zhao et al., 2019) use LSTM to learn the grammar and temporal features of stateful protocols. Specifically, they employ Long short-term memory (LSTM) as the encoder and decoder of the sequence-to-sequence (seq2seq) model (Sutskever et al., 2014). Seq2seq model is an encoder-decoder model structure that can handle input and output sequences of different lengths. The encoder LSTM model learns the features of the protocol via captured network traces, while the decoder LSTM model is used to generate fuzzing inputs. Passive network trace-based state machine learning methods are easy to operate and fast-running. However, the quality of the constructed state machine depends on the coverage of captured traffics. In practice, it is hard to capture a comprehensive set of message types and sequences, causing the constructed communication model to lack parts of uncaptured states or state transitions.
@@
Active learning (labeled as “TAAL” in Table LABEL:table-inputgen column 4) methods involve learning the communication model during the fuzzing process (de Ruiter and Poll, 2015; McMahon Stone et al., 2018; Daniel et al., 2018; Fiterau-Brostean et al., 2020; Aichernig et al., 2021; Fiterau-Brostean et al., 2022; Sagonas and Typaldos, 2023; Yu et al., 2023; Garbelini et al., 2022b; Pham et al., 2020; Luo et al., 2023; Garbelini et al., 2022a). These approaches can be categorized on the basis of whether the global state set is predefined. The first category does not define the global state set in advance, i.e., meaning it does not predetermine the number and nature of possible states in the state machine. This approach employs automata active learning algorithms to discern the state machine of the target. The learning algorithms are based on user-defined input/output alphabets and mappers between alphabets and concrete messages. Starting from an empty state machine, these algorithms iteratively propose and refine the model by interacting with the target protocol implementation, ceasing only when no counterexamples to the learned state machine are found. Most works in this category (de Ruiter and Poll, 2015; McMahon Stone et al., 2018; Daniel et al., 2018; Fiterau-Brostean et al., 2020; Aichernig et al., 2021; Fiterau-Brostean et al., 2022; Sagonas and Typaldos, 2023) utilize Angluin’s algorithm, defining input alphabets based on protocol specifications and translating these into actual messages using message templates. DYNPRE (Luo et al., 2024a) incorporates an adaptive message rewriting technique to manage session-specific identifiers while interacting with the server. It uses byte-level mutations and server feedback to deduce the meaning and format of messages, identify different message types, and create a precise and minimal protocol state machine based on observed message patterns. Conversely, the second category pre-defines the state set through a rule-based method to circumvent the complexities of automata learning algorithms, and learns the transitions between states by mutating known messages. For example, AFLNET (Pham et al., 2020) uses response message status codes to infer the current protocol state, mutating real message sequences to uncover transitions. Bleem (Luo et al., 2023) utilizes the Scapy library to parse messages and abstract them into various message types by retaining all fields of the enumeration type. This strategy is based on empirical observations from more than 50 protocols supported by Scapy, where different enumeration field values typically signify distinct packet or frame types. Bleem then uses these abstracted message traces to construct a guiding graph for fuzzing. Another example is Braktooth (Garbelini et al., 2022a), which defines eight rules that map messages to states based on message characteristics. It operates as a proxy between the PUT and a standard protocol stack, mutating communications to explore additional state transitions. Similarly, Garbelini et al.(Garbelini et al., 2022b) establish mapping rules to identify states and learn the state machine using capture traces (i.e., pcap files).
@@
Program-Analysis-Assisted Approaches (labeled as “PAL” in Table LABEL:table-inputgen column 4): Compared with traffic-analysis-based approaches, program-analysis-assisted approaches additionally use internal execution information to construct the communication model. In general, internal execution information includes the results of static and dynamic program analysis, which requires not only the access to the program but also the availability of analysis frameworks such as program instrumentation tools. Based on the type of internal execution information used, existing work can be divided into execution-trace-based and state-variable-based. Execution-Trace-Based approaches recognize different internal execution states according to the execution trace of the target. For example, ICS3Fuzzer(Fang et al., 2021) dynamically instruments the target supervisory software to collect the trace. By comparing the identity of the execution traces, ICS3Fuzzer can distinguish whether the PUT is in a different state. The state-variable-based approaches detect protocol state transitions by tracking the value changes of state variables during input processing (Natella, 2022; McMahon Stone et al., 2022; Ba et al., 2022; Qin et al., 2023). These approaches are based on the simple observation that most protocol implementations use certain variables to store the current state. Therefore, they identify these variables as state variables and use their values to distinguish between different states. For example, StateAFL (Natella, 2022) identifies possible state variables by identifying long-lived data structures in memory snapshots. Similarly, STATEINSPECTOR (McMahon Stone et al., 2022) identifies state variables by locating memory regions in heap memory that kept the same values in the execution of each message sequence. Differently, SGFuzz (Ba et al., 2022) identifies state variables through regular expressions by automatically extracting all enum type variables that are assigned at least once. The insight behind this approach is based on the investigation that most protocol implementations use enum-type state variables. STATELIFTER (Shi et al., 2023b) views the loops in protocol parsers as the core of the state machine. Through static analysis, it traverses the loop structures in the code and maps the different paths that each loop iteration may execute to distinct states. The dependencies between loop iterations are treated as state transitions within the state machine. ParDiff (Zheng et al., 2024) uses static symbolic analysis to extract finite state machines (FSMs) from protocol implementations by identifying and converting the relevant message parsing constraints into ordered state transitions.
Some works have also used program analysis-assisted approaches to construct the data models of protocol implementations (Caballero et al., 2007; Comparetti et al., 2009; Sun et al., 2023; Shi et al., 2023a). Polyglot (Caballero et al., 2007) leverages dynamic binary analysis to extract protocol information by closely monitoring how a program processes network data, revealing intricate protocol semantics and structures without source code dependency. Recent work Netlifter(Shi et al., 2023a) uses static analysis to derive precise protocol specifications directly from source code, employing Abstract Format Graphs (AFG) to capture and visualize complex data structure relationships and dependencies with high accuracy. Spenny(Sun et al., 2023) integrates dynamic analysis with symbolic execution to precisely reverse-engineer industrial control system protocols.
@@ 4.2. Task Scheduling
In the realm of recent protocol fuzzing research, the scheduling phase has been distinctly categorized based on the methodology employed to handle state-related complexities. This classification leads to two primary categories: Hierarchical Approaches and Monolithic Approaches.
Hierarchical Approaches decompose the scheduling process into two discrete phases: (1) Inter-state scheduling: This phase involves selecting a state to fuzz using a state scheduling algorithm based on the priority or relevance of the state. (2)Intra-state scheduling: Once the target state is selected, a general scheduling algorithm is applied to optimize fuzzing within that state. By separating these phases, hierarchical approaches allow for more nuanced control over the fuzzing process. For example, if we want to test a protocol after a handshake is completed, the inter-state scheduling phase will first prioritize this state. Then, the intra-state scheduling phase will generate specific test cases to be executed within the post-handshake state using traditional scheduling methods (e.g., seed scheduling, byte scheduling, mutation strategy scheduling). In this paradigm, the heuristics used by the scheduling process mainly fall into three categories, namely rarity-preferred (SRHS in Table LABEL:table-inputgen column 5), performance-preferred (SPHS in Table LABEL:table-inputgen column 5), and complexity-preferred (SCHS in Table LABEL:table-inputgen column 5), as detailed in Table 3. Rarity-preferred heuristics allocate more resources to seldomly exercised states, hypothesizing that these states harbor more undiscovered adjacent states or code logics (Pham et al., 2020; Natella, 2022; Sun et al., 2019; Borcherding et al., 2023). Performance-preferred heuristics prioritize states demonstrating higher code coverage or bug discovery rates (Chen et al., 2019; Pham et al., 2020; Natella, 2022; Garbelini et al., 2020, 2022c). Furthermore, some works utilize complexity-preferred heuristics, favoring states with greater complexity (i.e., connected to more basic blocks) or deeper states (i.e., further from the initial state) (Gascon et al., 2015; Fang et al., 2021). For example, ICS3Fuzzer (Fang et al., 2021) inclines to choose the deeper states and those states that exercise more basic blocks. As a generation-based fuzzer, Pulsar (Gascon et al., 2015) calculates the weight of all states that can be reached from the current state and then selects the state that has the maximum weight to be tested next. In detail, the weight of a state is calculated as the sum of all mutable fields in a fixed number of transitions. However, since all these state selection algorithms are implemented and evaluated separately on different platforms and targets, it is difficult to make a fair comparison and achieve conclusive findings. Liu et al. (Liu et al., 2022a) evaluate the three existing state selection algorithms of AFLNet (Pham et al., 2020), including a rarity-preferred algorithm, an algorithm that randomly selects states, and a sequential state selection algorithm. They find that these algorithms achieved very similar results in terms of code coverage. They attribute the reasons to the coarse-grained state abstraction of AFLNET and the inaccurate estimation of the state productivity. Therefore, they propose the AFLNETLEGION algorithm (Liu et al., 2022a) to address these issues, which is based on a variant of the Monte Carlo tree search algorithm (Liu et al., 2020).
In contrast, monolithic approaches employ a single, unified scheduling phases where state-related information is integrated directly into the scheduling algorithm. This means that the scheduler considers the performance of seeds in various states as part of its decision-making process, rather than treating state transitions and test case generation as separate steps. For example, SGFuzz (Ba et al., 2022) divides the states into rare states and normal states according to the exercised times. When assigning energy to seeds, it calculates the proportion of the rare states that are exercised by each seed and adds this proportion as one of the parameters on the basis of the original power scheduling algorithm. In a similar way, SGFuzz assigns more energy to the seeds containing state transitions that correspond to the expected protocol behaviors. This is because SGFuzz expects that these valid state transitions are easier to be mutated to other invalid state transitions, thus incurring error handling logic. Similarly, LTL-Fuzzer (Meng et al., 2021) also schedules the entire seed, prioritizing seeds that are closer to the target code locations during execution.
In addition to state-related information, many works utilize other categories of information for scheduling purposes. However, the scheduling algorithms based on these categories of information are generally universal and have been well discussed in the literature (Manes et al., 2021; Zhu et al., 2022). Therefore, we did not discuss them in detail.